OpenStudy (anonymous):

Someone Help ME 1.Solve P = 2(l + w) for l. What are the missing values in the table? 2.Create your own unique literal equation and solve for one of the variables. Show your work. Then, using complete sentences, explain how you solved for the variable you chose. 3.Using complete sentences, explain how solving a literal equation is similar to or different from simplifying an expression such as 6 - 2(52 + 7) ÷ 4. 4Using complete sentences, explain what might happen if the order of operations was used to solve a literal equation.

OpenStudy (anonymous):

P w l 14 2 ? 22 8 ? THIS IS THE TABLE FOR QUESTION ONE

jimthompson5910 (jim_thompson5910):

first let's solve for L (I'm using upper case because lower case L looks like a one) P = 2(L + W) P = 2L + 2W P - 2W = 2L 2L = P - 2W L = (P - 2W)/2

jimthompson5910 (jim_thompson5910):

to fill out that table, you just plug in the given values in each row so the first row has P = 14 and W = 2, so L = (P - 2W)/2 L = (14 - 2*2)/2 ... plug in P = 14 and W = 2 L = (14 - 4)/2 L = 10/2 L = 5

jimthompson5910 (jim_thompson5910):

so the L in the first row is 5

jimthompson5910 (jim_thompson5910):

I'll let you do the second row

OpenStudy (anonymous):

okay thanks

jimthompson5910 (jim_thompson5910):

you're welcome